# Neurons Speak in Ranges: Breaking Free from Discrete Neuronal Attribution

### v1

@@ 5 Activation Ranges-guided Intervention
Table 4: Evaluation of selected models on the AG-News dataset using neuron and range masking techniques. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance. For Bert and GPT-2 50% and for Llama-3.2-3B 30% neurons selected.
@@
5.1 Experimental Setup
Models. This study employs a range of transformer-based architectures, including encoder and decoder models such as fine-tuned BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2020), GPT-2 (Radford et al., 2019), and pretrained Llama-3.2-3B (Grattafiori, 2024). We incorporate our methodology at the penultimate layer, abiliation for layer selection is provided in the Appendix E. The training details for the models are provided in appendix B.
Datasets. We consider various classification based tasks; sentiment analysis (IMDB, Maas et al. (2011)), (SST2, Socher et al. (2013)), emotion detection (Dair-Ai/Emotions Saravia et al. (2018)), news classification (AG-News (Zhang et al., 2015)) and article content classification tasks (DBPedia-14 (Zhang et al., 2015)).
Metrics. For performance evaluation, we employ two quantitative metrics: prediction accuracy and the model’s confidence score for the ground truth. Initially, we established baseline measurements of both accuracy and confidence for all concepts $C$ in the unmodified model. Post-intervention measurements are recorded for both the target concept $c$ and auxiliary concepts $c^{\prime}$. The effectiveness and precision of neuron identification can be assessed through two key metrics: (1) the magnitude of performance degradation for concept $c$, and (2) the extent of unintended impact on auxiliary concepts $c^{\prime}$. Throughout our analysis, we denote the accuracy and confidence metrics for concept $c$ as Acc and Conf respectively, while corresponding measurements for auxiliary concepts $c^{\prime}$ are represented as CAcc and CConf.
5.2 Results and Analysis
Table 3 presents comprehensive results across five datasets, demonstrating the effectiveness of our range-based masking approach compared to traditional neuron masking. Range-based masking better preserves auxiliary concepts across all settings performance (CAcc, CConf), suggesting more precise concept isolation.
Multi-class classification tasks (AG-News, Emotions, DBPedia-14) demonstrate more pronounced intervention effects. Range masking achieves substantial primary task degradation while maintaining auxiliary concept performance, particularly evident in AG-News results (Table 4). The class-level analysis reveals asymmetric impacts: Classes 0 and 1 show larger performance drops (-0.919, -0.912 for GPT-2) compared to Classes 2 and 3 (-0.678, -0.887), suggesting varying degrees of concept distribution across neurons.
Model architecture and training methodology significantly influence intervention outcomes. BERT, trained directly on the classification tasks with a dedicated classification head, demonstrates remarkable resilience with minimal accuracy degradation (-0.015 on DBPedia-14) overall, under range masking it maintains auxiliary concept performance. This robustness likely stems from its task-specific fine-tuning, which enables more distributed and redundant concept representations. In contrast, GPT-2, despite being fine-tuned but trained in an autoregressive manner, shows substantially higher vulnerability with major drops in AG-NEWS ($\Delta_{acc}=-0.849$) and DBPedia-14 ($\Delta_{acc}=-0.782$). This increased sensitivity may be attributed to its autoregressive training objective, which potentially leads to more sequential and less redundant concept encodings. Llama-3.2-3B, evaluated in a few-shot setting without task-specific training, experiences the most severe degradation across all datasets (often exceeding -0.900), suggesting that pre-trained representations without task-specific fine-tuning are more vulnerable to targeted neuron interventions. These patterns indicate that task-specific training, particularly with classification-oriented objectives, may foster more robust and distributed concept representations that are more resilient to neuron interventions.
The effectiveness of range-based masking in preserving auxiliary concepts while achieving targeted performance degradation suggests that concept encoding in neural networks may follow activation range patterns rather than being strictly neuron-specific. This is particularly evident in the AG-News results, where range masking maintains better auxiliary concept scores (CAcc, CConf) compared to neuron masking across all classes and models.
Figure 6: Accuracy comparison between Neuronal Range manipulation (green) and complete neuron manipulation (orange) methods. Both approaches show similar primary task degradation (solid lines) but differ significantly in auxiliary task stability (dashed lines). The Neuronal Range method maintains more stable auxiliary task performance than Maximal Activation’s steep degradation after 50%. Model GPT-2
On binary classification tasks (IMDB, SST2), both masking approaches show moderate performance drops in targeted concepts, with BERT experiencing accuracy decreases of -0.184 and -0.060 respectively under range masking. The relatively modest impact on primary task performance, coupled with minimal auxiliary concept interference, indicates significant neuronal redundancy in binary classification tasks multiple neurons appear to encode similar sentiment-related features.
Percentage Masking Effect As the proportion of masked neurons increases, the performance improvement over the maximum activation baseline becomes more pronounced. Beyond a critical masking threshold, the baseline performance deteriorates exponentially, while our proposed method maintains stable performance up to 100% masking. This phenomenon can be explained by two key factors. First, as we ablate a higher percentage of neurons, the probability of encountering polysemantic neurons increases as a statistical consequence. Second, neurons identified through maximum activation are more likely to exhibit discriminatory behavior between concepts or demonstrate strong conceptual associations. when analyzing neurons with lower activation patterns (not strongly associated with concepts), discrete classification becomes an increasingly crude metric, as evident from Figure 6. After 50% masking, we see a steep drop in the general performance of the model, in the case of neuron masking. This suggests the need for a more nuanced approach to mapping information as to neural activations as blocking/manipulating a higher percentage of the model’s representation creates a significant deviation from the original model’s behavior. Our range-based methodology provides such precision, enabling more accurate characterization of neuronal behavior across the activation spectrum.
6 Related Work
While we have discussed closely related approaches in Section 2, here, we briefly review additional relevant techniques. Circuit discovery identifies groups of neurons that jointly encode concepts, providing a structured view of model behavior (Marks et al., 2024; Conmy et al., 2023; Olah et al., 2020). However, extracting circuits is computationally intensive and lacks fine-grained neuron-level attribution. Gradient-based methods attribute predictions to input features by tracking gradients through the network, with integrated gradients (Sundararajan et al., 2017; Dai et al., 2022) being a widely used approach. However, they struggle with polysemanticity, as they do not disentangle overlapping concepts within neurons. Causal analysis methods intervene on internal components to assess their role in encoding concepts. Causal tracing measures the effect of corrupting activations on model performance (Vig et al., 2020; Meng et al., 2022), while causal mediation analysis quantifies information propagation through neurons (Vig et al., 2020). Although effective, these techniques require costly perturbation experiments. Beyond neuron-level analysis, representation-level methods examine hidden states and their relationship to model outputs. Logit lens (nostalgebraist, 2020) and tuned lens (Belrose et al., 2023) project intermediate representations into the model’s vocabulary space to track information flow, while sparse probing (Gurnee et al., 2023) compresses hidden representations into sparse, interpretable subspaces. While prior work has advanced interpretability, most methods rely on discrete neuron-to-concept mappings, which fail to account for polysemanticity. Our work extends activation-based approaches by introducing activation ranges as the fundamental unit of interpretability to enable more precise concept attribution and intervention.
7 Conclusion
In this work, we challenged traditional assumptions about neuron interpretability by reframing polysemanticity as a resource rather than a limitation in interpreting neurons. Through an in-depth analysis of encoder- and decoder-based large language models across multiple text classification datasets, we uncovered that neuronal activations for individual concepts exhibit distinct, Gaussian-like distributions. This discovery allows for a more precise understanding of how neurons encode multiple concepts, enabling us to move beyond coarse, monolithic neuron-to-concept mappings. Building upon these insights, we proposed $\mathsf{NeuronLens}$, a novel range-based framework for neuron interpretation and manipulation. $\mathsf{NeuronLens}$ offers fine-grained control that reduces interference with unrelated concepts by attributing specific activation ranges within neurons to individual concepts. Extensive empirical evaluations demonstrated that $\mathsf{NeuronLens}$ outperforms existing methods in maintaining concept-specific precision while minimizing unintended side effects. The hyperparameter $\tau$ was set to ensure fair comparison across methods. Future work could explore mathematically deriving optimal $\tau$ values for individual neurons. Additionally, our work proved that the fundamental unit of interpretation can be a range within the activation space, which can be exploited for better interpretability.
8 Impact Statement
Appendix A Limitations
While $\mathsf{NeuronLens}$ can disentangle polysemanticity to a degree using Gaussian Like Distribution, it is unable to completely disentangle concepts encoded in the polysemantic neurons, because there still is a significant overlap in the distributions of concepts in activations. Additionally, in this work, we use $\tau$ to be a fixed value of 2.5 to make the comparison of approaches fair, but $\tau$ selection can be optimized to be more sophisticated. We also get results primarily from the penultimate layer, and not the intermediate or earlier layers, however, we do give ablation and rationale for this choice in Appendix E
Appendix B Training Details
Appendix C Saliency details
Max activations. Frankle & Carbin (2019) extract high neural activations as a saliency ranking metric relying upon the rationale that maximally activating neurons are salient as these neurons play a critical role in controlling the model’s output, highlighting their importance for a concept $c$.To identify them, the column-wise mean of absolute neuronal activations in $H^{l}_{c}$, $H^{l}_{c}$ is defined in Section2.3, is computed, given that high negative activations also carry significant signals (Voita et al., 2023). The magnitude of the means is then considered as a ranking for concept $c$.
Probe analysis. Dalvi et al. (2019b) train a linear classifier on the hidden representations $H^{l}_{c}$ to distinguish between concepts. The learned model weights are then utilized as a saliency ranking. This process involves learning a weight matrix $W\in\mathbb{R}^{d\times|c|}$, where $d$ is the hidden dimension and $|c|$ is the number of concept classes. The absolute weight values of each row in the weight matrix are used as a ranking for the importance of each neuron for a given concept. To prevent the emergence of redundant solutions characterized by minimal variations in the weights, the probe is trained using the elastic regularization technique.
Probeless. Antverg & Belinkov (2022) examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. The element-wise difference between mean vectors is computed as $r=\sum_{c,c^{\prime}\in C}|q(c)-q(c^{\prime})|$, where $r\in\mathbb{R}^{d}$ and $d$ is the hidden dimension. The final neuron saliency ranking is obtained by sorting $r$ in descending order.
Appendix D Full Results
@@
Appendix E Layer Ablation
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
Model Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf

### v2

@@ D.1 Dampening
Table 5: Evaluation of Llama-3.2-3B on a DBPedia-14 dataset using neuron and range masking techniques. 30% neurons were selected. Dampening factor used is $a=0.125$. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Neuron Masking and Activation Range Masking show deviations from the baseline performance. PPL $\Delta$ and MMLU $\Delta$ show changes in perplexity and MMLU scores, respectively.
Class Base Values Neuron Masking (Deviations) Activation Range Masking (Deviations)
Acc Conf CAcc CConf Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$ Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$
Class 0 1.000 0.576 1.000 0.563 -0.919 -0.545 -0.281 -0.309 3.161 -0.020 -0.924 -0.545 -0.276 -0.285 0.640 -0.010
Class 1 1.000 0.526 1.000 0.567 -0.988 -0.467 -0.246 -0.270 3.578 -0.015 -0.805 -0.466 -0.193 -0.206 0.725 0.015
Class 2 1.000 0.441 1.000 0.575 -0.864 -0.391 -0.461 -0.323 2.891 -0.030 -0.869 -0.389 -0.346 -0.282 0.718 0.005
Class 3 1.000 0.461 1.000 0.573 -0.974 -0.439 -0.411 -0.346 3.036 -0.025 -0.970 -0.438 -0.282 -0.283 0.653 0.010
Class 4 1.000 0.839 1.000 0.541 -0.382 -0.597 -0.367 -0.317 2.997 $0$0.382 -0.597 -0.334 -0.284 0.691 0.020
Class 5 1.000 0.339 1.000 0.568 -0.970 -0.326 -0.239 -0.246 3.503 0.010 -0.970 -0.325 -0.197 -0.187 0.810 0.015
Class 6 1.000 0.810 1.000 0.545 -0.233 -0.638 -0.194 -0.276 3.126 -0.010 -0.241 -0.637 -0.174 -0.203 0.697 -0.010
Class 7 1.000 0.595 1.000 0.562 -0.210 -0.382 -0.206 -0.226 3.037 $0$0.179 -0.376 -0.123 -0.143 0.546 0.015
Class 8 1.000 0.417 1.000 0.574 -0.310 -0.416 -0.335 -0.297 3.001 0.020 -0.346 -0.416 -0.200 -0.187 0.624 0.015
Class 9 1.000 0.526 1.000 0.567 -0.820 -0.465 -0.327 -0.264 3.369 -0.030 -0.809 -0.463 -0.213 -0.189 0.596 0.000
Class 10 1.000 0.505 1.000 0.569 -0.691 -0.466 -0.389 -0.314 3.732 $0$0.696 -0.465 -0.267 -0.198 0.695 -0.015
Class 11 1.000 0.497 1.000 0.569 -0.873 -0.432 -0.472 -0.289 3.070 -0.030 -0.865 -0.427 -0.335 -0.205 0.594 -0.015
Class 12 1.000 0.573 1.000 0.563 -0.720 -0.452 -0.295 -0.221 3.410 -0.045 -0.723 -0.451 -0.190 -0.163 0.595 0.035
Class 13 1.000 0.567 1.000 0.564 -0.951 -0.537 -0.226 -0.189 2.995 $0$0.955 -0.536 -0.157 -0.150 0.672 0.005
Table 5 presents a comparative analysis of two intervention strategies, neuron masking and activation range masking, when employing the Dampening technique with $\alpha=0.5$. The evaluation spans 14 classes and utilizes the metrics: accuracy (Acc), confidence (Conf), class-wise accuracy (CAcc), class-wise confidence (CConf), alterations in perplexity (PPL), and MMLU score.
A consistent trend emerges across the primary metrics (Acc, Conf, CAcc, and CConf), where activation range masking demonstrates superior performance over neuron masking. Interventions based on activation ranges lead to a notably smaller decline in the accuracy and confidence associated with auxiliary concepts. For example, in Class 3, while neuron masking results in an accuracy drop of -0.974 in the targeted class and auxiliary class accuracy decrease of -0.411, activation range masking, despite a comparable accuracy reduction in the targeted class (-0.970), shows a less severe impact on auxiliary class accuracy (-0.283). This pattern of activation range masking better preserves auxiliary class performance, is evident across all evaluated classes.
Examining the broader effects on language modeling capabilities reveals significant distinctions between the two approaches. Neuron masking results in a considerable rise in perplexity (PPL), with increases ranging from +2.891 to +3.732 across all the classes. Furthermore, it tends to cause more pronounced negative shifts in MMLU scores, reaching as low as -0.045. Conversely, activation range masking results in substantially smaller increments in perplexity, falling within the +0.546 to +0.810 range, and frequently results in improved or minimally altered MMLU scores, with gains up to +0.035.
D.2 Mean Replacement
Another approach of activation replacement discussed in the literature[Suau et al., 2021] is replacing it with the mean activation value. We provide the results for this type of replacement in Table 6.
The mean replacement strategy corresponds to setting the intervention function to $\phi(x)=\mu$, where $\mu$ is the mean activation of the neuron $x$ computed over a general next-token prediction task on the Wikipedia[Foundation, ].
Table 6: Evaluation of Llama-3.2-3B on a DBPedia-14 dataset using neuron and range masking techniques. 30% neurons were selected. Mean Activation $\mu$ is used as replacement value. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Neuron Masking and Activation Range Masking show deviations from the baseline performance. PPL $\Delta$ and MMLU $\Delta$ show changes in perplexity and MMLU scores, respectively.
Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$ Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$
Class 0 1.000 0.576 1.000 0.563 -1.000 -0.576 -0.685 -0.554 7.681 -0.025 -1.000 -0.576 -0.551 -0.545 0.687 -0.005
Class 1 1.000 0.526 1.000 0.567 -1.000 -0.526 -0.554 -0.550 8.437 -0.030 -1.000 -0.526 -0.356 -0.517 0.583 0.015
Class 2 1.000 0.441 1.000 0.575 -0.995 -0.441 -0.697 -0.556 7.567 -0.015 -0.995 -0.440 -0.574 -0.536 0.520 -0.010
Class 3 1.000 0.461 1.000 0.573 -1.000 -0.461 -0.766 -0.561 8.005 -0.015 -1.000 -0.461 -0.538 -0.534 0.543 0.010
Class 4 1.000 0.839 1.000 0.541 -1.000 -0.838 -0.724 -0.528 8.239 0.010 -0.995 -0.838 -0.502 -0.503 0.565 0.005
Class 5 1.000 0.339 1.000 0.568 -1.000 -0.339 -0.616 -0.551 7.753 0.010 -1.000 -0.339 -0.382 -0.510 0.552 0.005
Class 6 1.000 0.810 1.000 0.545 -0.313 -0.805 -0.549 -0.531 7.880 -0.005 -0.292 -0.805 -0.336 -0.499 0.547 0.020
Class 7 1.000 0.595 1.000 0.562 -1.000 -0.592 -0.491 -0.535 7.413 -0.010 -0.995 -0.591 -0.267 -0.449 0.462 0.000
Class 8 1.000 0.417 1.000 0.574 -0.928 -0.414 -0.632 -0.556 7.688 0.015 -0.934 -0.414 -0.298 -0.489 0.495 0.015
Class 9 1.000 0.526 1.000 0.567 -1.000 -0.526 -0.611 -0.544 8.057 -0.035 -1.000 -0.526 -0.370 -0.482 0.467 0.015
Class 10 1.000 0.505 1.000 0.569 -0.998 -0.505 -0.642 -0.558 8.791 -0.020 -0.998 -0.505 -0.406 -0.485 0.484 0.005
Class 11 1.000 0.497 1.000 0.569 -1.000 -0.497 -0.719 -0.543 7.903 0.025 -1.000 -0.497 -0.447 -0.459 0.397 -0.005
Class 12 1.000 0.573 1.000 0.563 -0.904 -0.572 -0.629 -0.543 8.046 -0.005 -0.896 -0.571 -0.375 -0.484 0.425 0.000
Class 13 1.000 0.567 1.000 0.564 -1.000 -0.566 -0.526 -0.533 7.543 -0.025 -0.998 -0.566 -0.341 -0.481 0.464 -0.010
In Table 6, we assess the effect of mean replacement using both neuron masking and activation range masking. In every class, neuron masking results in more severe degradation than range-based masking across all auxiliary and general metrics.
Across metrics (Acc, Conf, CAcc, and CConf), activation range masking consistently outperforms neuron masking. The degradation in accuracy and confidence of auxiliary concepts is significantly lower under range-based interventions. For instance, in Class 3, neuron masking causes a drop of -1.000 in Acc and -0.766 in CAcc, whereas activation range masking yields a similar Acc drop (-1.000) but a substantially smaller decline in CAcc (-0.538). A similar pattern repeats across all classes; for example, in Class 0, neuron masking results in CAcc of -0.685 while activation range masking yields -0.551. In Class 7, neuron masking shows a CAcc of -0.491 compared to -0.267 for activation range masking.
Beyond auxiliary class performance, we observe substantial differences in how the two masking methods affect general language modelling capabilities. Neuron masking leads to a large increase in perplexity (PPL), ranging from +7.413 to +8.791 which is catastrophic, across classes, and induces more negative shifts in MMLU scores (as low as -0.035 for Class 9, and also for Class 0 with -0.025, Class 1 with -0.030, Class 10 with -0.020, and Class 13 with -0.025). In contrast, activation range masking results in substantially smaller increases in perplexity (+0.397 to +0.687) and often yields improved or near-zero changes in MMLU scores (up to +0.020 for Class 6, and several positive values like +0.015 for Class 1, Class 8, and Class 9).
D.3 Adaptive Dampening
We propose a novel replacement method in which the intervention function $\phi(x)$ applies runtime-controlled dampening based on the distance of the observed activation $x$ from the center of a predefined activation range. Specifically, the dampening factor $a(x)$ is linearly scaled according to the distance of $x$ from the mean $\mu$ of the neuron’s activation distribution, within the range $[\mu-2.5\sigma,\mu+2.5\sigma]$.
Let $\beta\in[0,1]$ denote the maximum dampening factor applied at the range boundaries. Then:
$a(x)=\beta\cdot\frac{|x-\mu|}{2.5\sigma},\quad\text{and}\quad\phi(x)=a(x)\cdot x.$
This ensures that when $x=\mu$ (the center of the range), $a(x)=0$ and the activation is fully suppressed via $\phi(x)=0$. At the boundaries ($x=\mu\pm 2.5\sigma$), $a(x)=\beta$, and the activation is minimally dampened. Values within the range are scaled proportionally based on their normalized distance from the mean. This adaptive dampening mechanism suppresses values near the mean while preserving those closer to the range edges.
The dampening factor $\beta$ can be optimized for different neurons based on the concept information that neuron provides. For this work, we use $\beta=0.5$ across all neurons.
Table 7: Evaluation of Llama-3.2-3B on a DBPedia-14 dataset using neuron and range masking techniques. 30% neurons were selected. Adaptive Dampening factor used is $\beta=0.5$. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Neuron Masking and Activation Range Masking show deviations from the baseline performance. PPL $\Delta$ and MMLU $\Delta$ show changes in perplexity and MMLU scores, respectively.
Class Base Values Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf PPL $\Delta$ MMLU $\Delta$
Class 0 1.000 0.576 1.000 0.563 -0.927 -0.543 -0.215 -0.217 0.487 -0.015
Class 1 1.000 0.526 1.000 0.567 -0.791 -0.451 -0.134 -0.109 0.543 0.000
Class 2 1.000 0.441 1.000 0.575 -0.828 -0.380 -0.277 -0.215 0.540 -0.010
Class 3 1.000 0.461 1.000 0.573 -0.958 -0.432 -0.230 -0.214 0.492 0.010
Class 4 1.000 0.839 1.000 0.541 -0.346 -0.579 -0.261 -0.218 0.521 0.015
Class 5 1.000 0.339 1.000 0.568 -0.960 -0.319 -0.140 -0.116 0.609 -0.015
Class 6 1.000 0.810 1.000 0.545 -0.236 -0.613 -0.130 -0.122 0.524 -0.010
Class 7 1.000 0.595 1.000 0.562 -0.243 -0.388 -0.108 -0.080 0.408 0.005
Class 8 1.000 0.417 1.000 0.574 -0.440 -0.414 -0.152 -0.088 0.465 0.030
Class 9 1.000 0.526 1.000 0.567 -0.799 -0.459 -0.182 -0.131 0.445 0.005
Class 10 1.000 0.505 1.000 0.569 -0.684 -0.451 -0.222 -0.130 0.513 -0.010
Class 11 1.000 0.497 1.000 0.569 -0.836 -0.420 -0.308 -0.155 0.440 -0.005
Class 12 1.000 0.573 1.000 0.563 -0.720 -0.451 -0.172 -0.095 0.444 0.025
Class 13 1.000 0.567 1.000 0.564 -0.941 -0.530 -0.142 -0.098 0.502 0.010
In Table 7 we evaluate the adaptive dampening variant of the replacement function. This approach outperforms both neuron masking and static activation masking across all metrics.
In auxiliary class metrics, adaptive dampening yields much smaller degradation. Auxilary class accuracy (CAcc) and confidence (CConf) show significantly reduced drops compared to other methods. For example, in Class 0, CAcc drops only $-0.215$ compared to $-0.685$ under neuron masking and $-0.551$ under hard activation masking. The effect is consistent across classes, with most CAcc and CConf drops staying well below $-0.3$, and in many cases below $-0.15$.
Language modeling metrics show this approach to be exceptionally efficient. Perplexity increases are minimal, remaining within $+0.408$ to $+0.609$, substantially lower than all hard-masking variants. MMLU deltas also stay close to zero, with several classes showing improvement (e.g., Class 8: $+0.030$, Class 4: $+0.015$). Notably, no class suffers significant MMLU degradation.
Appendix E Training Details
In the case of Bert-based models, record the activation of the CLS token, In the case of GPT-2 and Llama models, we record the last token output when the class token is being predicted. The intervention is applied to the appropriate token on the residual stream.
Appendix F Compute Details
All experiments, including activation extraction and interventions on large language models (LLMs), were conducted using an NVIDIA RTX 3090 GPU equipped with 24GB of VRAM. 64GB RAM.
Appendix G Saliency details
Max activations. Frankle and Carbin [2019] extract high neural activations as a saliency ranking metric relying upon the rationale that maximally activating neurons are salient as these neurons play a critical role in controlling the model’s output, highlighting their importance for a concept $c$.To identify them, the column-wise mean of absolute neuronal activations in $H^{l}_{c}$, $H^{l}_{c}$ is defined in Section2.3, is computed, given that high negative activations also carry significant signals [Voita et al., 2023]. The magnitude of the means is then considered as a ranking for concept $c$.
Probe analysis. Dalvi et al. [2019b] train a linear classifier on the hidden representations $H^{l}_{c}$ to distinguish between concepts. The learned model weights are then utilized as a saliency ranking. This process involves learning a weight matrix $W\in\mathbb{R}^{d\times|c|}$, where $d$ is the hidden dimension and $|c|$ is the number of concept classes. The absolute weight values of each row in the weight matrix are used as a ranking for the importance of each neuron for a given concept. To prevent the emergence of redundant solutions characterized by minimal variations in the weights, the probe is trained using the elastic regularization technique.
Probeless. Antverg and Belinkov [2022] examine individual neurons, without the need for auxiliary classifiers, using the element-wise difference between mean vectors. The element-wise difference between mean vectors is computed as $r=\sum_{c,c^{\prime}\in C}|q(c)-q(c^{\prime})|$, where $r\in\mathbb{R}^{d}$ and $d$ is the hidden dimension. The final neuron saliency ranking is obtained by sorting $r$ in descending order.
Table 8: Performance drops relative to Baseline configuration (i.e.: unaltered model’s performance) for three techniques: Probeless, Probe, and Max. All values show the difference from Base values. Results are for Emotions dataset on the GPT-2 model using 30% salient neurons of each method. Metrics are detailed in 2.1.
Probeless Probe Max
Class Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
Appendix H Hyperparameter Ablation
For target concept, $\tau$ values $0.3-2.4$ show decreasing accuracy/confidence, stabilizing at $\tau=2.4$ (accuracy $0.6126$). Beyond $2.4$, negligible additional degradation occurs, indicating we’ve captured the complete target concept activation range. Importantly, while target performance stabilizes after $\tau=2.4$, auxiliary task performance declines after $\tau=2.7$. Complement accuracy stays above 0.93 until then before dropping to 0.8795 at $\tau=4.5$. This aligns with normal distribution properties where 95-99% of values fall within ±2.5 standard deviations.
Table 9: Performance metrics for varying $\tau$ values.
$\tau$ Acc Conf CAcc CConf
0.3 0.9021 0.8858 0.9452 0.9358
0.6 0.8439 0.8185 0.9424 0.9327
0.9 0.7801 0.7486 0.9391 0.9263
1.2 0.7295 0.6950 0.9340 0.9174
1.5 0.6834 0.6482 0.9337 0.9093
1.8 0.6424 0.6141 0.9331 0.9000
2.1 0.6184 0.5926 0.9327 0.8910
2.4 0.6126 0.5858 0.9314 0.8846
2.7 0.6024 0.5798 0.9280 0.8800
3.0 0.5971 0.5776 0.9234 0.8777
3.3 0.5963 0.5786 0.9173 0.8753
3.6 0.5970 0.5794 0.9097 0.8729
3.9 0.5976 0.5802 0.9020 0.8698
4.2 0.5967 0.5798 0.8908 0.8642
4.5 0.5967 0.5798 0.8795 0.8577
Appendix I Layer Ablation
I.1 Statistical Results
We analyze concept level activation distributions across all 12 layers of GPT-2, measuring kurtosis (where a value of 3.0 indicates a Gaussian distribution), skewness (where 0 indicates symmetry), and practical normality in Table 10:
Table 10: Statistical analysis of different layers showing skewness, kurtosis, and Kolmogorov- Smirnov test results. GPT2 model. AG-News Dataset
Layer Kurtosis Skewness Practical Normality(10%)
1 3.9314 0.0430 0.7913
2 3.7622 -0.0091 0.9525
3 3.4109 -0.0143 0.9870
4 3.5582 -0.0073 0.9801
5 3.6145 0.0051 0.9730
6 3.5318 0.0086 0.9769
7 3.3461 0.0083 0.9880
8 3.2763 0.0037 0.9870
9 3.2267 0.0039 0.9860
10 3.2057 0.0029 0.9899
11 3.2105 -0.0002 0.9912
12 3.2061 -0.0014 0.9919
These results show that kurtosis values converge toward 3.0 (the Gaussian ideal) as layers progress, skewness values remain near zero across all layers, and practical normality scores are close to 1 across all layers. Importantly, if the activations were not clustered into continuous intervals and were in disconnected islands of activations, these would be reflected in the score for the practical normality and other statistical metrics.
I.2 Qualitative Results
We expanded our visualization approach shown in Figures figs. 8, 8, 10, 10, 12, 12, 14, 14, 16, 16, 18 and 18 in Figure 4) to all layers in the model. The visualizations demonstrate an interesting progression: while all layers exhibit Gaussian-like distributions on the class level, class concepts aren’t separated in the activation spectrum Gaussians of the early layers. This aligns with the understanding that lower layers capture more basic features rather than high-level semantic features like class. However, distinct concept-level Gaussian distributions begin forming as early as layers 5-6, becoming increasingly separable in deeper layers.
Figure 7: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 1
Figure 8: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 2
Figure 9: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 3
Figure 10: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 4
Figure 11: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 5
Figure 12: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 6
Figure 13: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 7
Figure 14: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 8
Figure 15: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 9
Figure 16: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 10
Figure 17: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 11
Figure 18: Neuronal Activation Patterns of six neurons on AG-News dataset. Layer 12
I.3 Masking Results
Layer Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
Layer Class Base Values Neuron Masking Activation Range Masking
Acc Conf CAcc CConf Acc Conf CAcc CConf Acc Conf CAcc CConf
Appendix J Class Wise Results
Table 13: Evaluation of selected models on the AG-News dataset using neuron and range masking techniques. Acc represents class accuracy, Conf denotes class prediction probability, and CAcc and CConf refer to average accuracy and average class prediction probability across other classes, respectively. The Base Values indicate the baseline model performance, while Activation Range Masking and Neuron Masking show deviations from the baseline performance. For GPT-2 50% and for Llama-3.2-3B 30% neurons selected.